how to calculate magnetic quantum number? explain clearly with an example

Grade:12

3 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Student

The magnetic quantum number has values between -l and +l. When l =1, for example, m can have three values: -1, 0, and +1. Because you know  that the subshell designation for l =1 is "p", you now know that the p orbital has three components  px, py, and pz. Notice how the subscripts are related to a three-dimensional coordinate system, x, y, and z. The chart below shows a summary of the quantum numbers:

 

Principal Quantum Number (n) Azimuthal Quantum Number (l) Subshell Designation Magnetic Quantum Number (m) Number of orbitals in subshell
1 0 1s 0 1
2 0
1
2s
2p
0
-1 0 +1
1
3
3 0
1
2
3s
3p
3d
0
-1 0 +1
-2 -1 0 +1 +2
1
3
5
4 0
1
2
3
4s
4p
4d
4f
0
-1 0 +1
-2 -1 0 +1 +2
-3 -2 -1 0 +1 +2+3
1
3
5
7

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

 

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

To describe the magnetic quantum number m you begin with an atomic electron's angular momentum, L, which is related to its quantum number l\, by the following equation:

\mathbf{L} = \hbar\sqrt{l(l+1)}

where \hbar = h/2\pi is the reduced planck constant.

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

Aakash Dutta
13 years ago

this is very simple